OpenStudy (yanasidlinskiy):

There are 15 members of the show choir. In how many ways can you arrange 4 members in the front row? Permutation or combination? Solve. @johnweldon1993

OpenStudy (anonymous):

In permutations, the order of items matters. In combination, order doesn't matter. Since here, order does matter, you'd do 15P4 or P(15,4), equivalent to 15*14*13*12.

OpenStudy (yanasidlinskiy):

What. Ok. Are you sure it is Permutations?

OpenStudy (anonymous):

Yes. If the four people in the front row are called Bob, Bill, Stacy, and Emily, and they were arranged that way, it would be different than Emily, Bob, Stacy, and Bill.

OpenStudy (yanasidlinskiy):

How come it's not combinations?

OpenStudy (anonymous):

In combinations, the arrangement of items does not matter.

OpenStudy (anonymous):

So Bob, Bill, Stacy, and Emily would be the same as Emily, Bob, Stacy, and Bill if it were combinations.

OpenStudy (anonymous):

But, the question asks how many different ways can they be arranged, which tells us that it;s a permutation.

OpenStudy (anonymous):

I believe the answer is 32760, that is 15*14*13*12

OpenStudy (anonymous):

Number of ways = 15 P 4 = 32760 ways

OpenStudy (anonymous):

The above is true

OpenStudy (yanasidlinskiy):

@david111 I don;'t want the answer. I need step by step.

OpenStudy (anonymous):

15*14*13*12

OpenStudy (yanasidlinskiy):

@johnweldon1993 Would it be like this??\[_{15}P _{4}\]

OpenStudy (anonymous):

So you're selecting 4 from 15 to be arranged in the front row. That's 15P4, which is equal to 15*14*13*12, which gives you 32760.

OpenStudy (johnweldon1993):

Indeed @YanaSidlinskiy Everything that has been said so far has been true :)

OpenStudy (yanasidlinskiy):

Ok. Tell me if im right. Ok?

OpenStudy (anonymous):

Go ahead.

OpenStudy (yanasidlinskiy):

Wait... I'm kind of stuck... This is what I've got so far. \[_{15}P _{4} = \frac{ 15! }{ (15-4) }\] Don't know what to do next...

OpenStudy (anonymous):

It's 15!/(15-4)!, and n! = n*(n-1)*(n-2) and so on until *1. So for example, 6! would be 6*5*4*3*2*1, and 7! would be 7*6*5*4*3*2*1. So here you have 15!/11!.

OpenStudy (anonymous):

\[\frac{ 15! }{ 11! }\]

OpenStudy (yanasidlinskiy):

What? I don't get this>_<

OpenStudy (johnweldon1993):

Well now you evaluate look at it like this \[\large \frac{15!}{(15 - 4)!} = \frac{15 \times 14 \times 13\times12\times11\times10\times9\times8\times7\times6\times5\times4\times3\times2}{11\times10\times9\times8\times7\times6\times5\times4\times3\times2\times1}\] Now look that after 11....they all just cancel out...so we just have \[\large \frac{15!}{(15 - 4)!} = \frac{15 \times 14 \times 13\times12\cancel{\times11\times10\times9\times8\times7\times6\times5\times4\times3\times2}}{\cancel{11\times10\times9\times8\times7\times6\times5\times4\times3\times2\times1}}\] So all we have is \[\large 15 \times 14 \times 13 \times 12 = \space?\]

OpenStudy (anonymous):

And since 15! can divide by 11! because 15! is 15*14*13*12*11*10*9*8*7*6*5*4*3*2*1, you can just cancel out the 11*10*9*8*7*6*5*4*3*2*1 so you have a numerator of 15*14*13*12 and a denominator of 1 because the 11! has canceled out.

OpenStudy (anonymous):

Lol a repeat of john's answer.

OpenStudy (yanasidlinskiy):

That's what I was looking for @johnweldon1993 That's what I actually needed!:) Thanx:) For everyone who helped!:)

OpenStudy (anonymous):

Kudos to Mr. John and you're welcome. :)

OpenStudy (johnweldon1993):

Haha you did all the work @aurorablue01 so enjoy the earned medal... and anytime @YanaSidlinskiy :)

OpenStudy (anonymous):

Lol, thanks :)

OpenStudy (anonymous):

Now, I've gotta go do me some math homework, so see you later, peeps!

OpenStudy (anonymous):

The order doesn't matter it's combination